OpenStudy (anonymous):

Leave the answer in your lowest form 5 9/16-3 3/4

OpenStudy (gw2011):

5 9/16 = 5 9/16 = 4 25/16 -3 3/4 = -3 12/16 = -3 12/16 Therefore, 4 25/16 - 3 12/16 = 1 13/16

OpenStudy (anonymous):

superstar the 5 is whole number and 9/16 is fraction -3 whole number 3/4 as fraction and i have to show my workings

OpenStudy (gw2011):

The answer is 1 and 13/16. Since the fraction of 9/16 associated with the 5 is smaller than the fraction of 12/16 associated with the -3, you basically borrow 1 from the 5 and this reduces the 5 to 4 and the borrowed 1 is 16/16 which you add to the 9/16 and you get 25/16.